We are able to solve this problem using guess & check, but are unable to write it as an equation. The ages of John & Mary total 27 years. Mary's age plus twice John's age is 40. How old is each person?
Found 2 solutions by Alan3354, scott8148:
Answer by Alan3354(69443)

The ages of John & Mary total 27 years. Mary's age plus twice John's age is 40. How old is each person?
-------------------
M+ J = 27
M+2J = 40
------------- subtract
-J = -13
J = 13
M = 27

Answer by scott8148(6628)

allow me to translate from English to Algebra ...

"The ages of John & Mary total 27 years" ___ j + m = 27

"Mary's age plus twice John's age is 40" ___ m + 2j = 40

subtracting the 1st eqn from the 2nd ___ j = 13

substituting ___ 13 + m = 27 ___ m = 14